The sum of ages of 4 family members is 68. Four years ago the sum was 53. How old is the youngest family member? Please EXPLAIN!
Found 2 solutions by idkweirdo, ankor@dixie-net.com:
Answer by idkweirdo(15)

The youngest family is 13 years old. Because if u divide 68 years of age by the four members of the family, the youngest has to be at least 17 years old. Four years ago, the youngest member was was 13.
LEave comments!!
Good luck

Answer by ankor@dixie-net.com(22740)

The sum of ages of 4 family members is 68. Four years ago the sum was 53. How old is the youngest family member? Please EXPLAIN!
:
let their ages (going from the youngest) = a, b, c ,d
:
a + b + c + d = 68
:
the 4 year ago equation
(a-4) + (b-4) + (c-4) + (d - 4) = 53
a + b + c + d - 16 = 53
then
a + b + c + d = 53 + 16
a + b + c + d = 69; which is not possible,
:
We can assume that "a" was not born 4 years ago
Then 4 yrs ago:
(b-4) + (c-4) + (d-4) = 53
b + c + d - 12 = 53
b + c + d = 53 + 12
b + c + d = 65
:
Therefore:
a = 68 - 65
a = 3 yrs old